• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name johna

    Log entry time 00:01:29 on April29,2011

    Entry number 357482

    keyword=Swing shift summary

    Shift started with hall in controlled access while MCC finished beam studies. We were informed at ~17:30 that MCC was ready to start beam delivery, so we ended controlled access.

    Beam delivery started at ~18:00, but after checking raster and doing harp scan, there was difficulty providing high current beam. By 20:00, they had steady beam at 90uA, and we started taking data while they continuted to try and acheive high current.

    From ~18:20 on, had 118uA beam, although worse trip problems than usual. Hall C stayed at 165 through the shift.

    3He production running: runs 3983-3986, extimate about 8k good events total (about 2k/hr when running well)